# A progress on the binary Goldbach conjecture

### v8

@@ 2.1. Application of Density of Points to Partitions
Let $\mathbb{H}\subset\mathbb{N}$ such that $\mathcal{D}(\mathbb{H})>\frac{1}{2}$. Then every sufficiently large $n\in\mathbb{N}$ has representation of the form
@@
Appealing to Proposition 2.3 we can write
@@
By the uniqueness of the axes of CoPs we can write
#$\#\left\{\mathbb{L}_{[x],[y]}~{}\hat{\in}~{}\mathcal{C}(n)|~{}\{x,y\}\cap\mathbb{H}\neq\emptyset\right\}$ $=\nu(n,\mathbb{H})+\#\left\{\mathbb{L}_{[x],[y]}~{}\hat{\in}~{}\mathcal{C}(n)|~{}x\in\mathbb{H},~{}y\in\mathbb{N}\setminus\mathbb{H}\right\}.$
Let us assume $\nu(n,\mathbb{H})=0$ then it follows by appealing to Definition 2.2 and according to the proof of Proposition 2.3
@@
Let :=$\mathbb{R}:=\left\{m\in\mathbb{N}|~{}\mu(m)\neq 0\right\}$. Then every sufficiently large $n\in\mathbb{N}$ can be written in the form
@@
where $\mu(z_{1})=\mu(z_{2})\neq 0$.
@@
#$\#\left\{\mathbb{L}_{[x],[y]}~{}\hat{\in}~{}\mathcal{C}(n)|~{}\{x,y\}\cap\mathbb{R}\neq\emptyset\right\}$ $=\nu(n,\mathbb{R})+\#\left\{\mathbb{L}_{[x],[y]}~{}\hat{\in}~{}\mathcal{C}(n)|~{}x\in\mathbb{R},~{}y\in\mathbb{N}\setminus\mathbb{R}\right\}.$
Let us assume $\nu(n,\mathbb{R})=0$ then it follows by appealing to Definition 2.2 and Theorem 2.6
@@
One could ever hope and dream of this strategy to work when we replace the set $\mathbb{R}$ of square-free integers with the set of prime numbers. There we would certainly ran into complete deadlock, since the prime in accordance with the prime number theorem have density zero. Any success in this regard could conceivably work by introducing some exotic forms of the notion of density of points and carefully choosing a subset of the integers that is somewhat dense among the set of integers and covers the primes. We propose a strategy somewhat akin to the above method for possibly getting a handle on the binary Goldbach conjecture and it’s variants. Before that we state and prove a conditional theorem concerning the binary Goldbach conjecture.
@@
Let $\mathbb{B}\subset\mathbb{N}$ such that $\mathbb{P}\cap\mathbb{N}_{n}\subset||\mathcal{C}(n,\mathbb{B})||$ with
@@
where $\eta(n)=|\mathcal{C}(n,\mathbb{B})|$. Then $\nu(n,\mathbb{P})>0$ for all sufficiently large values of $n\in 2\mathbb{N}$.
@@
First let us upper and lower bound the density of points in the CoP $\mathcal{C}(n,\mathbb{B})$ with weight belonging to the set of the primes $\mathbb{P}$ so that we obtain the inequality
@@
for all sufficiently large values of $n$. Appealing to the uniqueness of the axes of CoPs, we can write
#$\#\left\{\mathbb{L}_{[x],[y]}~{}\hat{\in}~{}\mathcal{C}(n,\mathbb{B})|~{}\{x,y\}\cap\mathbb{P}\neq\emptyset\right\}$ $=\nu(n,\mathbb{P})+\#\left\{\mathbb{L}_{[x],[y]}~{}\hat{\in}~{}\mathcal{C}(n,\mathbb{B})|~{}x\in\mathbb{P},~{}y\in\mathbb{B}\setminus\mathbb{P}\right\}.$
Let us assume to the contrary $\nu(n,\mathbb{P})=0$, then it follows that no two points in the CoP $\mathcal{C}(n,\mathbb{B})$ with weight in the set $\mathbb{P}$ are axes partners, so that under the requirement
@@
2.2. A Strategy to Prove the Binary Goldbach Conjecture by Circles of Partition
In this subsection we propose series of steps that could be taken to confirm the truth of the binary Goldbach conjecture. We enumerate the strategies chronologically as follows:
  1. $\bullet$
    First construct a CoP $\mathcal{C}(n,\mathbb{B})$ such that $\mathbb{P}\cap\mathbb{N}_{n}\subset||\mathcal{C}(n,\mathbb{B})||$ and that
@@ 1. $\bullet$
  2. $\bullet$
    Next we remark that the following inequality also holds and this can be obtain by replacing the weight set $||\mathcal{C}(n)||$ with the set $||\mathcal{C}(n,\mathbb{B})||$.
@@ 2. $\bullet$
  3. $\bullet$
    Appealing to the uniqueness of the axes of CoPs, we can write
    #$\#\left\{\mathbb{L}_{[x],[y]}~{}\hat{\in}~{}\mathcal{C}(n,\mathbb{B})|~{}\{x,y\}\cap\mathbb{P}\neq\emptyset\right\}$
@@ 3. $\bullet$
  4. $\bullet$
    Let us assume $\nu(n,\mathbb{P})=0$ then it follows by appealing to Definition 2.2 and in the sense of the proof of Proposition 2.3
@@ 4. $\bullet$
3. The asymptotic squeeze principle and the Binary Goldbach Conjecture - An asymptotic proof of the binary Goldbach conjecture
@@
In our paper [10], we introduced and developed the method of circles of partition. This method is underpinned by a combinatorial structure that encodes certain additive properties of the subsets of the integers and invariably equipped with a certain geometric structure that allows to view the elements as points in the plane whose weights are just elements of the underlying subset. We call this combinatorial structure the circle of partition and is refereed to as the set of points
@@
Each point in this set - except the center point - must have a uniquely distinct point that is joined by a line which we refer to as an axis of the CoP. We denote an axis of a CoP with $\mathbb{L}_{[x],[y]}$ and an axis contained in the CoP as
$\mathbb{L}_{[x],[y]}~{}\hat{\in}~{}\mathcal{C}(n,\mathbb{M})\mbox{ which means}[x],[y]\in\mathcal{C}(n,\mathbb{M})\mbox{ with }x+y=n.$
The method of circles of partition and their associated structures have been well advanced in [12], where the corresponding points have complex numbers as their weights and a line (axis) joining co-axis points. The following structure was considered as a complex circle of partition
@@
with $\mathbb{M}\subseteq\mathbb{N}$. We abbreviate this complex additive structure as cCoP. The condition $\Im(z)^{2}=\Re(z)(n-\Re(z)$ is referred to as the circle condition and it guarantees that all points on the cCoP lie on a circle in the complex. This circle is the embedding circle of the cCoP $\mathcal{C}^{o}(n,\mathbb{C}_{\mathbb{M}})$, denoted as $\mathfrak{C}_{n}$. The embedding circles of cCoPs have the property that they reside fully inside those embedding circles with a relatively larger generators, except the origin as a common point [12]. For each axis, we make the following assignment
$\mathbb{L}_{[z_{1}],[z_{1}]}~{}\hat{\in}~{}\mathcal{C}(n,\mathbb{C}_{\mathbb{M}})\mbox{ which means }[z_{1}],[z_{2}]\in\mathcal{C}(n,\mathbb{C}_{\mathbb{M}})\mbox{ with }z_{1}+z_{2}=n.$
The structure of the complex circle of partition is much more versatile and has extra structures that are not readily available in the theory of circle of partition. Most notably, for each axis $\mathbb{L}_{[z],[n-z]}$ of a cCoP there exists
@@
a conjugate axis, where $\overline{[z]},\overline{[n-z]}$ denotes the corresponding conjugate points. The space occupied by an embedding circle of partition and correspondingly outside the embedding circle turns out to be very interesting, as this can utilised in studying a certain ordering principle of the points of two interacting axes of distinct cCoPs. Much more striking is the fact, which is a natural consequence of the circle condition, that
@@
for any axis $\mathbb{L}_{[z_{1}],[z_{2}]}\in\mathcal{C}^{o}(n,\mathbb{C}_{\mathbb{M}})=\{[z]\mid z,n-z\in\mathbb{C}_{M},\Im(z)^{2}=\Re(z)\left(n-\Re(z)\right)\}$. The squeeze principle [11] can be considered as a black box for studying the binary Goldbach conjecture and its variant. A slightly different version of this principle appears in [12]. For the sake of the reader, we provide a brief recap of this elegant principle as below
Remark 3.1.
@@
$\Re(v_{1})<\Re(v_{2})\mbox{ and }\Re(w_{1})<\Re(w_{2})$ (3.1)
then there exists an axis $\mathbb{L}_{[\Re(v_{2})],[\Re(w_{1})]}~{}\hat{\in}~{}\mathcal{C}(n+s,\mathbb{B})$ with $0<s<t$. Hence the cCoP $\mathcal{C}^{o}(n+s,\mathbb{C}_{\mathbb{B}})$ not empty and so is the cCoP $\mathcal{C}^{o}(n+s,\mathbb{C}_{\mathbb{M}})$.
@@
From the existence of an axis $\mathbb{L}_{[v_{1}],[w_{1}]}~{}\hat{\in}~{}\mathcal{C}^{o}(n,\mathbb{C}_{\mathbb{M}})$ follows $\Re(w_{1})=n-\Re(v_{1})$. With the requirement (4.2) we get
$\Re(w_{1})>n-\Re(v_{2}).$ (3.2)
On the other hand from the existence of an axis $\mathbb{L}_{[v_{2}],[w_{2}]}~{}\hat{\in}~{}\mathcal{C}^{o}(n+t,\mathbb{C}_{\mathbb{M}})$ follows $\Re(w_{2})=n+t-\Re(v_{2})$ and with the requirement (4.2) and the result (3.2) we get
@@
By virtue of the requirements $w_{1},v_{2}\in\mathbb{C}_{\mathbb{B}}$ and $n+s=\Re(w_{1})+\Re(v_{2})$ there is an axis $\mathbb{L}_{[\Re(v_{2})],[\Re(w_{1})]}~{}\hat{\in}~{}\mathcal{C}(n+s,\mathbb{B})$ and hence holds $\mathcal{C}^{o}(n+s,\mathbb{C}_{\mathbb{B}})\neq\emptyset$. Since $\mathbb{B}\subset\mathbb{M}$, it follows immediately that $\mathbb{C}_{\mathbb{B}}\subset\mathbb{C}_{\mathbb{M}}$ and therefore $\mathcal{C}^{o}(n+s,\mathbb{C}_{\mathbb{M}})\neq\emptyset$. This completes the proof. ∎
@@
The Lemma 4.2 referred to as the squeeze principle may be regarded as a fundamental tool set for investigating the viability of dividing integers of a particular parity, utilizing constituent elements originating from a specific subset of the integers. The mechanism operates by discerning a pair of cCoPs that are both non-vacuous and share a common base set. Subsequently, supplementary cCoPs that are non-vacuous and have generators restrained within the interstice of these two generators are identified. This principle may be applied in a resourceful manner to investigate the overarching matter of the practicality of divvying up numbers such that each addend is a member of the identical subset of positive integers.
@@ 3.1. Application to the Binary Goldbach Conjecture
$\mathbb{L}_{[z_{1}],[z_{2}]}~{}\hat{\in}~{}\mathcal{C}^{o}(n,\mathbb{C}_{\mathbb{N}})$ and $\mathbb{L}_{[w_{1}],[w_{2}]}~{}\hat{\in}~{}\mathcal{C}^{o}(n+t,\mathbb{C}_{\mathbb{N}})$ for $\Re(z_{1})<\Re(z_{2})$ and $\Re(w_{1})<\Re(w_{2})$ such that $\Re(z_{1})\neq\Re(w_{1})$ and $\Re(z_{2})\neq\Re(w_{2})$ with $z_{2}\in\mathbb{C}_{\mathbb{P}}$ and $w_{1}\in\mathbb{C}_{\mathbb{P}}$ for $t\geq 4$.
@@
Let us choose a prime number $\Re(w_{1})\leq\frac{n+t}{2}$ and choose a prime number $\Re(z_{2})\in(\frac{n}{2},n)$ for all $n\geq 10$ with $n\equiv 0\pmod{2}$, which is feasible by virtue of the prime number theorem. If $\Re(z_{1})\neq\Re(w_{1})$ and $\Re(z_{2})\neq\Re(w_{2})$ then there is nothing to do. Without loss of generality, suppose that $\Re(z_{1})=\Re(w_{1})$ then obviously $\Re(z_{2})\neq\Re(w_{2})$ since $n+t>n$. We note that $\pi(\frac{n+t}{2})\geq 3$ for all $n\geq 10$ with $t\geq 4$ so that we can choose a prime number $\Re(w^{{}^{\prime}}_{1})\leq\frac{n+t}{2}$ such that $\Re(w^{{}^{\prime}}_{1})\neq\Re(w_{1})$. Thus we replace $\Re(w_{1})$ with $\Re(w^{{}^{\prime}}_{1})$ and obtain the axes $\mathbb{L}_{[z_{1}],[z_{2}]}~{}\hat{\in}~{}\mathcal{C}^{o}(n,\mathbb{C}_{\mathbb{N}})$ and $\mathbb{L}_{[w^{{}^{\prime}}_{1}],[w^{{}^{\prime}}_{2}]}~{}\hat{\in}~{}\mathcal{C}^{o}(n+t,\mathbb{C}_{\mathbb{N}})$ for $\Re(z_{1})<\Re(z_{2})$ and $\Re(w^{{}^{\prime}}_{1})<\Re(w^{{}^{\prime}}_{2})$ such that $\Re(z_{1})\neq\Re(w^{\prime}_{1})$. If $\Re(z_{2})\neq\Re(w^{{}^{\prime}}_{2})$ then we are done; otherwise, we choose another prime number $\Re(w^{{}^{\prime\prime}}_{1})$ such that $\Re(w^{{}^{\prime\prime}}_{1})\neq\Re(w^{{}^{\prime}}_{1})$ and $\Re(w^{{}^{\prime\prime}}_{1})\neq\Re(w_{1})$ since $\pi(\frac{n+t}{2})\geq 3$ for all $n\geq 10$ and $t\geq 4$. By virtue of our construction, we obtain finally the axes of cCoPs $\mathbb{L}_{[z_{1}],[z_{2}]}~{}\hat{\in}~{}\mathcal{C}^{o}(n,\mathbb{C}_{\mathbb{N}})$ and $\mathbb{L}_{[w^{{}^{\prime\prime}}_{1}],[w^{{}^{\prime\prime}}_{2}]}~{}\hat{\in}~{}\mathcal{C}^{o}(n+t,\mathbb{C}_{\mathbb{N}})$ for $\Re(z_{1})<\Re(z_{2})$ and $\Re(w^{{}^{\prime\prime}}_{1})<\Re(w^{{}^{\prime\prime}}_{2})$ such that $\Re(z_{1})\neq\Re(w^{{}^{\prime\prime}}_{1})$ and $\Re(z_{2})\neq\Re(w^{{}^{\prime\prime}}_{2})$ with $z_{2}\in\mathbb{C}_{\mathbb{P}}$ and $w^{{}^{\prime\prime}}_{1}\in\mathbb{C}_{\mathbb{P}}$ for $t\geq 4$. ∎
@@
(The prime number theorem ) Let $\pi(n)$ denotes the number of prime numbers no more than $n$. Then we have
@@
For $k\geq 89693$ there exists a prime number in the interval
@@
Let $p_{n}$ denotes the $n^{th}$ prime number, then
@@
There exist an $n_{o}\in\mathbb{N}$ such that for all even $n\geq n_{o}$ there exist an axis
$\mathbb{L}_{[z_{1}],[z_{2}]}~{}\hat{\in}~{}\mathcal{C}^{o}(n,\mathbb{C}_{\mathbb{N}})$ and $\mathbb{L}_{[w_{1}],[w_{2}]}~{}\hat{\in}~{}\mathcal{C}^{o}(n+t,\mathbb{C}_{\mathbb{N}})$ for $\Re(z_{1})\lesssim\Re(z_{2})$ and $\Re(w_{1})\lesssim\Re(w_{2})$ such that $\Re(z_{1})\lesssim\Re(w_{1})$ and $\Re(z_{2})\lesssim\Re(w_{2})$ with $z_{2}\in\mathbb{C}_{\mathbb{P}}$ and $w_{1}\in\mathbb{C}_{\mathbb{P}}$ for $t\geq 4$.
@@
Let us set $\Re(z_{2})$ to be a prime number and choose $\Re(z_{2})$ to be the $\pi(\frac{3n}{4})^{th}$ prime number. We note that via the prime number theorem holds
@@
We also note via basic power series identities, we can write
@@
Then with Lemma 3.8 and Lemma 3.6 we obtain
@@
$=\log n-\log\log n+O(1)$ (3.3)
It follows from (3.3), we can write for the product

### v9

@@ 3.1. Application of the density of Points to Partitions
Let $\mathbb{H}\subset\mathbb{N}$. If $\mathcal{D}(\mathbb{H})>\frac{1}{2}$, then every sufficiently large $n\in\mathbb{N}$ has a representation of the form
@@
Applying Proposition 3.3, we can write
@@
By the uniqueness of the axes of CoPs, we can write
$\#\left\{\mathbb{L}_{[x],[y]}~\hat{\in}~\mathcal{C}(n)~|~\{x,y\}\cap\mathbb{H}\neq\emptyset\right\}$ $=\nu(n,\mathbb{H})+\#\left\{\mathbb{L}_{[x],[y]}~\hat{\in}~\mathcal{C}(n)|~x\in\mathbb{H},~y\in\mathbb{N}\setminus\mathbb{H}\right\}.$
Let us assume that $\nu(n,\mathbb{H})=0$. This implies that each axes $\mathbb{L}_{[x],[y]}$ in this set must have the property that $x\in\mathbb{H}$ and $y\not\in\mathbb{H}$ or vise-versa. According to the proof of Proposition 3.3, we deduce
@@
Let $\mathbb{R}:=\left\{m\in\mathbb{N}~|~\mu(m)\neq 0\right\}$. Every sufficiently large $n\in\mathbb{N}$ can be written in the form
@@
where $\mu(z_{1})=\mu(z_{2})\neq 0$, where $\mu(\cdot)$ denotes the mobious function..
@@
$\#\left\{\mathbb{L}_{[x],[y]}~\hat{\in}~\mathcal{C}(n)|~\{x,y\}\cap\mathbb{R}\neq\emptyset\right\}$ $=\nu(n,\mathbb{R})+\#\left\{\mathbb{L}_{[x],[y]}~\hat{\in}~\mathcal{C}(n)|~x\in\mathbb{R},~y\in\mathbb{N}\setminus\mathbb{R}\right\}.$
Let us assume that $\nu(n,\mathbb{R})=0$. It from Theorem 3.6
@@
One could ever hope and dream of this strategy to work when we replace the set $\mathbb{R}$ of square-free integers with the set of prime numbers. We may have some difficulty because prime numbers have a zero density, in accordance with the prime number theorem. A progress along these paths could conceivably manifest by introducing some exotic forms of the notion of density of points and carefully choosing a subset of integers that is somewhat dense among the set of integers and contains prime numbers. Here, we propose an approach similar to the above method to possibly approach the binary Goldbach conjecture and it’s variants.
@@
[…]
@@
where $\eta(n)=|\mathcal{C}(n,\mathbb{B})|$. We have $\nu(n,\mathbb{P})>0$ for all sufficiently large values of $n\in 2\mathbb{N}$.
@@
We find an upper and a lower bound for the density of points in the CoP $\mathcal{C}(n,\mathbb{B})$ with weight belonging to the set of prime numbers $\mathbb{P}$. We obtain the inequality
@@
for all sufficiently large values of $n$. Exploiting the uniqueness of the axes of CoPs, we can write
$\#\left\{\mathbb{L}_{[x],[y]}~\hat{\in}~\mathcal{C}(n,\mathbb{B})~|~\{x,y\}\cap\mathbb{P}\neq\emptyset\right\}$ $=\nu(n,\mathbb{P})+\#\left\{\mathbb{L}_{[x],[y]}~\hat{\in}~\mathcal{C}(n,\mathbb{B})|~x\in\mathbb{P},~y\in\mathbb{B}\setminus\mathbb{P}\right\}.$
Let us assume that $\nu(n,\mathbb{P})=0$. This implies that no two points in the CoP $\mathcal{C}(n,\mathbb{B})$ with weight in the set $\mathbb{P}$ can be axes partners. Under the density requirement
@@
3.2. A strategy to prove the binary Goldbach conjecture using circles of partition
Here, we propose a series of steps that could be taken to confirm the truth of the binary Goldbach conjecture. We enumerate the strategies chronologically as follows:
  1. (1)
[…]
@@ 3.2. A strategy to prove the binary Goldbach conjecture using circles of partition
  2. (2)
    We remark that the following inequality also holds and this can be obtain by replacing the weight set $\mathcal{C}(n)||$ with the weight set $\mathcal{C}(n,\mathbb{B})||$:
@@
  3. (3)
    Exploiting the uniqueness of the axes of CoPs, we can write
    $\#\left\{\mathbb{L}_{[x],[y]}~\hat{\in}~\mathcal{C}(n,\mathbb{B})|~\{x,y\}\cap\mathbb{P}\neq\emptyset\right\}$
@@
  4. (4)
    We assume that $\nu(n,\mathbb{P})=0$ and deduce, in the sense of the proof of Proposition 3.3, the inequality
@@
4. The asymptotic squeeze principle and the Binary Goldbach Conjecture : An asymptotic proof of the binary Goldbach conjecture
@@
In section 2, we introduced and developed the method of circles of partition. This method is underpinned by a combinatorial structure that encodes certain additive properties of the subsets of the integers and invariably equipped with a certain geometric structure that allows to view the elements as points in the plane whose weights are elements of the underlying subset. We call this combinatorial structure the circle of partition and is refereed to as the set of points
@@
Each point in this set - except the center point - must have a uniquely distinct point that is joined by a line which we refer to as an axis of the CoP. We denote an axis of a CoP by $\mathbb{L}_{[x],[y]}$ and an axis contained in the CoP by
$\mathbb{L}_{[x],[y]}~\hat{\in}~\mathcal{C}(n,\mathbb{M})$
which means
$x],[y]\in\mathcal{C}(n,\mathbb{M})\quad\text{with}\quad x+y=n.$
In this section, we extend the base set of circles of partitions to the complex plane plane. Consequently, the weight of the axes points are now complex numbers.
Definition 4.1.
We denote the circle of partition with complex base set $\mathbb{C}_{\mathbb{M}}$ by
@@
with $\mathbb{M}\subseteq\mathbb{N}$. We call this complex additive structure a complex circle of partition and denote in short form by cCoP. The condition $\Im(z)^{2}=\Re(z)(n-\Re(z)$ is referred to as the circle condition and it guarantees that all points on the cCoP lie on a circle in the complex. This circle is the embedding circle of the cCoP $\mathcal{C}^{o}(n,\mathbb{C}_{\mathbb{M}})$, denoted by $\mathfrak{C}_{n}$. The embedding circles of cCoPs have the property that they reside fully inside those embedding circles with a relatively larger generators, except the origin as a common point. For each axis, we make the following assignment
$\mathbb{L}_{[z_{1}],[z_{1}]}~\hat{\in}~\mathcal{C}(n,\mathbb{C}_{\mathbb{M}})$
which means
$z_{1}],[z_{2}]\in\mathcal{C}(n,\mathbb{C}_{\mathbb{M}})\quad\text{with}\quad z_{1}+z_{2}=n.$
The structure of the complex circle of partition is more versatile and has extra features that are not readily available in the theory of circle of partition. Most notably, for each axis $\mathbb{L}_{[z],[n-z]}$ of a cCoP there exists
@@
a conjugate axis, where $\overline{[z]},\overline{[n-z]}$ denotes the corresponding conjugate points. The space occupied by an embedding circle of partition and, correspondingly, outside the embedding circle turns out to be very interesting. This statistic can be used to study a certain ordering principle of the weight of points of two interacting axes of distinct cCoPs. Much more striking is the fact, which is a natural consequence of the circle condition, that
@@
for any axis $\mathbb{L}_{[z_{1}],[z_{2}]}\in\mathcal{C}^{o}(n,\mathbb{C}_{\mathbb{M}})=\{[z]\mid z,n-z\in\mathbb{C}_{M},\Im(z)^{2}=\Re(z)\left(n-\Re(z)\right)\}$.
In the following, we develop the squeeze principle, a machinery that can be considered a black box for studying the binary Goldbach conjecture and its variant. Now, we make the following deductions:
@@
$\Re(v_{1})<\Re(v_{2})\quad\text{and}\quad\Re(w_{1})<\Re(w_{2})$
then there exists an axis $\mathbb{L}_{[\Re(v_{2})],[\Re(w_{1})]}~\hat{\in}~\mathcal{C}(n+s,\mathbb{B})$ with $0<s<t$. Hence, the cCoP $\mathcal{C}^{o}(n+s,\mathbb{C}_{\mathbb{B}})$ not empty and so is the cCoP $\mathcal{C}^{o}(n+s,\mathbb{C}_{\mathbb{M}})$.
@@
From the existence of an axis $\mathbb{L}_{[v_{1}],[w_{1}]}~\hat{\in}~\mathcal{C}^{o}(n,\mathbb{C}_{\mathbb{M}})$, we get $\Re(w_{1})=n-\Re(v_{1})$. With the requirement
$\Re(v_{1})<\Re(v_{2})\quad\text{and}\quad\Re(w_{1})<\Re(w_{2})$
we get
$\Re(w_{1})>n-\Re(v_{2}).$
On the other hand, by the existence of an axis $\mathbb{L}_{[v_{2}],[w_{2}]}~\hat{\in}~\mathcal{C}^{o}(n+t,\mathbb{C}_{\mathbb{M}})$, we get $\Re(w_{2})=n+t-\Re(v_{2})$ and with the requirement
$\Re(v_{1})<\Re(v_{2})\quad\text{and}\quad\Re(w_{1})<\Re(w_{2})$
and the inequality $\Re(w_{1})>n-\Re(v_{2})$, we get
@@
By the requirements $w_{1},v_{2}\in\mathbb{C}_{\mathbb{B}}$ and $n+s=\Re(w_{1})+\Re(v_{2})$, there is an axis $\mathbb{L}_{[\Re(v_{2})],[\Re(w_{1})]}~\hat{\in}~\mathcal{C}(n+s,\mathbb{B})$. Therefore $\mathcal{C}^{o}(n+s,\mathbb{C}_{\mathbb{B}})\neq\emptyset$. Since $\mathbb{B}\subset\mathbb{M}$, it follows immediately that $\mathbb{C}_{\mathbb{B}}\subset\mathbb{C}_{\mathbb{M}}$ and therefore $\mathcal{C}^{o}(n+s,\mathbb{C}_{\mathbb{M}})\neq\emptyset$. ∎
@@
The lemma 4.2 referred to as the squeeze principle may be regarded as a fundamental tool set for investigating the viability of dividing integers of a particular parity, utilizing constituent elements originating from a specific subset of the integers. The mechanism operates by discerning a pair of cCoPs that are both non-vacuous (non-empty) and share a common base set. Subsequently, supplementary cCoPs that are non-vacuous and have generators restrained within the interstice of these two generators are identified. This principle may be applied in a resourceful manner to investigate the overarching issue of the practicality of divvying up numbers such that each addend is a member of the identical subset of positive integers.
@@ 4.1. Application to the Binary Goldbach Conjecture
$\mathbb{L}_{[z_{1}],[z_{2}]}~\hat{\in}~\mathcal{C}^{o}(n,\mathbb{C}_{\mathbb{N}})$
and
$\mathbb{L}_{[w_{1}],[w_{2}]}~\hat{\in}~\mathcal{C}^{o}(n+t,\mathbb{C}_{\mathbb{N}})$
for $\Re(z_{1})<\Re(z_{2})$ and $\Re(w_{1})<\Re(w_{2})$ such that $\Re(z_{1})\neq\Re(w_{1})$ and $\Re(z_{2})\neq\Re(w_{2})$ with $z_{2}\in\mathbb{C}_{\mathbb{P}}$ and $w_{1}\in\mathbb{C}_{\mathbb{P}}$ for $t\geq 4$.
@@
We choose a prime number $\Re(w_{1})\leq\frac{n+t}{2}$ and choose a prime number $\Re(z_{2})\in(\frac{n}{2},n)$ for all $n\geq 10$ with $n\equiv 0\pmod{2}$, which is feasible by virtue of the prime number theorem. If $\Re(z_{1})\neq\Re(w_{1})$ and $\Re(z_{2})\neq\Re(w_{2})$, then there is nothing to do. Without loss of generality, suppose that $\Re(z_{1})=\Re(w_{1})$, then obviously $\Re(z_{2})\neq\Re(w_{2})$ since $n+t>n$. We note that $\pi(\frac{n+t}{2})\geq 3$ for all $n\geq 10$ with $t\geq 4$ so that we can choose a prime number $\Re(w^{{}^{\prime}}_{1})\leq\frac{n+t}{2}$ such that $\Re(w^{{}^{\prime}}_{1})\neq\Re(w_{1})$. Thus, we replace $\Re(w_{1})$ with $\Re(w^{{}^{\prime}}_{1})$ and obtain the axes $\mathbb{L}_{[z_{1}],[z_{2}]}~\hat{\in}~\mathcal{C}^{o}(n,\mathbb{C}_{\mathbb{N}})$ and $\mathbb{L}_{[w^{{}^{\prime}}_{1}],[w^{{}^{\prime}}_{2}]}~\hat{\in}~\mathcal{C}^{o}(n+t,\mathbb{C}_{\mathbb{N}})$ for $\Re(z_{1})<\Re(z_{2})$ and $\Re(w^{{}^{\prime}}_{1})<\Re(w^{{}^{\prime}}_{2})$ such that $\Re(z_{1})\neq\Re(w^{\prime}_{1})$. If $\Re(z_{2})\neq\Re(w^{{}^{\prime}}_{2})$, then we are done; otherwise, we choose another prime number $\Re(w^{{}^{\prime\prime}}_{1})$ such that $\Re(w^{{}^{\prime\prime}}_{1})\neq\Re(w^{{}^{\prime}}_{1})$ and $\Re(w^{{}^{\prime\prime}}_{1})\neq\Re(w_{1})$ since $\pi(\frac{n+t}{2})\geq 3$ for all $n\geq 10$ and $t\geq 4$. By virtue of our construction, we finally obtain the axes of cCoPs $\mathbb{L}_{[z_{1}],[z_{2}]}~\hat{\in}~\mathcal{C}^{o}(n,\mathbb{C}_{\mathbb{N}})$ and $\mathbb{L}_{[w^{{}^{\prime\prime}}_{1}],[w^{{}^{\prime\prime}}_{2}]}~\hat{\in}~\mathcal{C}^{o}(n+t,\mathbb{C}_{\mathbb{N}})$ for $\Re(z_{1})<\Re(z_{2})$ and $\Re(w^{{}^{\prime\prime}}_{1})<\Re(w^{{}^{\prime\prime}}_{2})$ such that $\Re(z_{1})\neq\Re(w^{{}^{\prime\prime}}_{1})$ and $\Re(z_{2})\neq\Re(w^{{}^{\prime\prime}}_{2})$ with $z_{2}\in\mathbb{C}_{\mathbb{P}}$ and $w^{{}^{\prime\prime}}_{1}\in\mathbb{C}_{\mathbb{P}}$ for $t\geq 4$. ∎
@@
(The prime number theorem ) Let $\pi(n)$ denote the number of prime numbers no more than $n$. We have
@@
For all $k\geq 89693$, there exists a prime number in the interval
@@
Let $p_{n}$ denotes the $n^{th}$ prime number. We have
@@
There exist an $n_{0}\in\mathbb{N}$ such that for all even numbers $n\geq n_{0}$ there exist an axis
$\mathbb{L}_{[z_{1}],[z_{2}]}~\hat{\in}~\mathcal{C}^{o}(n,\mathbb{C}_{\mathbb{N}})$
and
$\mathbb{L}_{[w_{1}],[w_{2}]}~\hat{\in}~\mathcal{C}^{o}(n+t,\mathbb{C}_{\mathbb{N}})$
with $\Re(z_{1})\lesssim\Re(z_{2})$ and $\Re(w_{1})\lesssim\Re(w_{2})$ such that $\Re(z_{1})\lesssim\Re(w_{1})$ and $\Re(z_{2})\lesssim\Re(w_{2})$ with $z_{2}\in\mathbb{C}_{\mathbb{P}}$ and $w_{1}\in\mathbb{C}_{\mathbb{P}}$ for $t\geq 4$.
@@
Let us set $\Re(z_{2})$ to be a prime number and choose $\Re(z_{2})$ to be the $\pi(\frac{3n}{4})^{th}$ prime number. We note that via the prime number theorem
@@
We also note that, via basic power series identities, we can write
@@
By the lemma 4.8 and the lemma 4.6, we obtain
@@
(4.1) $=\log n-\log\log n+O(1)$
From (4.1), we can write for the product